OpenStudy (anonymous):

the hamburger shop sells 500 sandwiches each day. They sell 100 more hamburgers than they do chicken sandwiches

OpenStudy (anonymous):

h + c = 500 c = h + 100

OpenStudy (anonymous):

i still have to solve it ?

OpenStudy (anonymous):

i guess... lets work it out together

OpenStudy (anonymous):

okay thanks how do we start

OpenStudy (anonymous):

do you want to use substitution or elimination?

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

ok do you know how to do that?

OpenStudy (anonymous):

no not really i need to solve the equation you gave me !

OpenStudy (anonymous):

omg sry doin a test brb... use the c= equation in the 1st one h+ h+100 = 500 and simplify